Your grandmother needs your help. She has $50,000 to invest. Part of this money is to be invested in noninsured bonds paying 15% annual interest. The rest of this money is to be invested in a government-insured certificate of deposit paying 7% annual interest. She told you that she requires $6000 per year in extra income from both of these investments. How much money should be placed in each investment?
Answer by macston(5194)

.
F=amount at 15%; S=amount at 7%
.
F+S=$50000
S=$50000-F Use this to substitute for S.
.
0.07S+0.15F=$6000 Substitute for S from above.
0.07($50000-F)+0.15F=$6000
$3500-0.07F+0.15F=$6000
0.08F=$2500
F=$31250
ANSWER 1: $31250 should be invested at 15%.
S=$50000-F=$50000-$31250=$18750
ANSWER 2: $18750 should be invested at 7%.
.
CHECK:
0.07S+0.15F=$6000
0.07($18750)+0.15($31250)=$6000
$1312.50+$4687.50=$6000
$6000=$6000